OpenStudy (anonymous):

find the equation of the line passing through the given points (0,1) (3,-8) pls help me

OpenStudy (anonymous):

use y=mx+b

OpenStudy (anonymous):

thanks going to try it now

OpenStudy (anonymous):

@ timo86m that formula you gave me was wrong i used the m=y1-y2/x2 and my answer is m=2/3 do you know about that formula?